OpenStudy (anonymous):

Solve x^2 + 8x + 8 = 0 by completing the square. Which equation is used in the process? a) (x + 4)^2= 0 b) (x + 4)^2= 4 c) (x + 4)^2= 8

OpenStudy (nurali):

x^2 + 8x + 8 = 0 x(x+8)+8=0 (x+4)^2-8=0 dividing by 8 1/8(x+4)^2=1 (x+4)^2=8

OpenStudy (anonymous):

Thank you! @Nurali

OpenStudy (nurali):

Anytime.